OpenStudy (anonymous):

just a general question. How do i find the one sided limit as x approaches 0 algebraically?

OpenStudy (anonymous):

Trace the graph /curve from the side of interest

OpenStudy (amistre64):

plug in numbers closer and closer to zero and see what they tend towards; algebraically there are deltas and epsilons to contend with

OpenStudy (amistre64):

\[\lim_{b->a}\frac{f(b)-f(a)}{b-a}\]

OpenStudy (amistre64):

or some such definition

OpenStudy (anonymous):

:( im still lost... here's the question \[\lim_{x \rightarrow -2+} (x/x+1)(2x+5)/(x^2+ x)\] supposd to find the limit algebraically. There is no graph.

OpenStudy (anonymous):

18/(6)=3

OpenStudy (anonymous):

1

OpenStudy (anonymous):

\[\lim_{x \rightarrow -2+} (x/x+1)(2x+5)/(x^2+ x) = 1\]

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%5Clim_ {x+%5Crightarrow+-2%2B}+%28x%2Fx%2B1%29%282x%2B5%29%2F%28x^2%2B+x%29